# Bidirectional Generative Pre-training for Improving Time Series Representation Learning

### v1

@@ 4.2 Visualization of Different Pre-training Strategies
Table 2: Compare BiTimelyGPT against various baselines for downstream seq2vec tasks. Best values are highlighted in bold.
@@
BiTimelyGPT 90.41 93.02 87.47 95.36 26.11 2.81 8.42 4.23
TimelyGPT 89.21 92.79 86.52 95.87 26.17 2.78 8.53 4.26
TimelyGPT w/o Conv 85.32 90.45 84.51 93.26 27.32 3.03 8.92 4.55
TimelyBERT 82.17 85.64 81.03 84.52 30.96 4.22 12.86 4.82
CRT 90.12 91.05 87.81 94.56 26.52 2.96 9.02 4.48
PatchTST 89.57 91.27 83.42 95.23 26.08 2.89 9.46 4.45
AutoFormer 78.86 84.21 78.51 88.56 32.18 4.13 13.29 4.95
FedFormer 76.43 81.65 75.53 85.21 31.11 4.36 13.82 4.75
TimesNet 83.58 85.96 79.38 89.26 29.95 4.19 13.65 4.83
TST 88.83 88.02 81.86 94.16 26.81 3.47 12.63 4.95
GPT 84.25 88.65 83.11 92.58 28.62 3.71 11.09 4.64
GPT-BAAR 86.17 89.73 85.09 93.71 28.17 3.60 10.62 4.66
TS2Vec 86.21 88.27 82.65 93.77 27.89 3.53 11.56 4.60
TS-TCC 86.06 89.74 84.66 93.25 29.32 4.09 13.64 4.86
TF-C 86.56 87.52 82.71 93.83 28.52 4.38 14.15 4.87
LSTM 80.15 76.51 78.83 86.95 30.15 4.95 14.37 5.05
LSTM-BAAR 82.22 81.05 82.30 87.96 28.06 4.53 11.78 4.82
BiLSTM 81.63 78.32 80.72 86.84 29.43 4.71 13.50 4.97
ELMo 82.15 81.60 80.98 88.52 28.15 4.66 12.42 4.76
4.3 Visualization of Bidirectionality
We investigated the effectiveness of bidirectionality in attention matrices on a downstream classification task. Through a case study in Figure 5, we compared the attention matrices from fine-tuned models of BiTimelyGPT, TimelyGPT (Song et al., 2023), GPT-2 (Radford et al., 2019), and TimelyBERT. For BiTimelyGPT  (Figure 5.b), we combined its lower and upper triangular retention matrices, $\overrightarrow{{\bm{R}}}^{L-1}$ and $\overleftarrow{{\bm{R}}}^{L}$, to demonstrate its bidirectional representations. Our analysis showed that BiTimelyGPT exhibited localized and concentrated attention on the key discriminative region, efficiently integrating bidirectional contexts. In contrast, TimelyGPT (Figure 5.c) exhibit more limited focus on the discriminative segment. GPT-2 (Figure 5.d) had a broader but less precise range of attention around these key features. TimelyBERT (Figure 5.e), despite having decay mechanisms in both directions, struggled to concentrate attention on the discriminative regions. These findings highlight the effectiveness of BiTimelyGPT in identifying discriminative patterns through its modeling of bidirectional contextualized representations.
Figure 6: Normalized cumulative singular values of the attention matrices. A more uniformly distributed curve of singular values, indicates a higher rank of the corresponding attention matrix.
4.4 Visualization of Attention Matrix Rankedness
We assessed the expressiveness of the attention matrices in various bidirectional transformers, including BiTimelyGPT, PatchTST, CRT, and TimelyBERT. We applied singular value decomposition (SVD) to the attention matrices across different layers and heads, visualizing the normalized cumulative singular values in Figure 6. Our analysis revealed that BiTimelyGPT conferred a more uniform distribution of singular values, effectively overcoming the low-rank bottleneck. In contrast, the singular value curves for other bidirectional baselines skewed towards the upper left corner, suggesting that a few largest singular values can account for most information in the attention matrices. Notably, TimelyBERT had the most long-tail spectrum distribution, due to its retention matrix $\overleftrightarrow{{\bm{R}}}=\hat{\bm{Q}}\hat{\bm{K}}^{\top}\odot\overleftrightarrow{{\bm{D}}}$ lacking softmax’s non-linearity. As a result, BiTimelyGPT demonstrated more expressive representation power, breaking the low-rank bottleneck common in other bidirectional transformers.
4.5 Performance on seq2vec Fine-Tuning Tasks
Baselines. For the classification and regression supervised fine-tuning tasks, we compared BiTimelyGPT against a broad range of baselines, including transformer PTMs, consistency-based PTMs, and recurrent models. Among transformer models, we included TimelyGPT, TimelyBERT, CRT, PatchTST, AutoFormer (Wu et al., 2021), FedFormer (Zhou et al., 2022), TST (Zerveas et al., 2021), and TimesNet (Wu et al., 2023). Consistency-based PTMs such as TS-TCC (Eldele et al., 2021), TS2Vec (Yue et al., 2022), and TF-C (Zhang et al., 2022a) were also evaluated; these models focus on learning effective representations by optimizing the distance between representations of positive and negative pairs via contrastive learning (Ma et al., 2023). We also assessed recurrent models like LSTM, BiLSTM, and ELMo. Additionally, to explore the efficacy of our proposed BAAR framework, we applied it to other autoregressive models GPT-2 and LSTM. We describe the architectures and parameters for BiTimelyGPT and baselines in Section B.2.
@@ 4.5 Performance on seq2vec Fine-Tuning Tasks
Results. As shown in Table 2, BiTimelyGPT demonstrated the best performance in the classification on Sleep-EDF, Epilepsy and EMG datasets and the regression on HR and SpO2 datasets. These results highlight the effectiveness of BiTimelyGPT’s bidirectionality, surpassing unidirectional TimelyGPT enhanced with convolution. Additionally, the data preservation during pre-training and full-rank retention matrices contributed considerably to BiTimelyGPT’s superior performance over other bidirectional transformers. CRT, employing a dropping-based pre-training, performed consistently well across datasets and achieved top results in the PTB-XL dataset. In contrast, masking-based PTMs typically lagged behind due to the distribution shift. Among these masking-based PTMs, Autoformer, Fedformer, and TimesNet, which rely heavily on time decomposition and frequency-domain information, were especially affected by distribution shift. This suggests the advantage of dropping-based pre-training regime in avoiding distribution shift over the masking-based pre-training methods. Furthermore, both consistency-based PTMs and recurrent models showed inferior performance in these tasks. The applications of the BAAR framework to GPT and LSTM resulted in improved performance, indicating its broad adaptability and effectiveness in enhancing various autoregressive models.
@@
Table 3: Ablation results of BiTimelyGPT w/o specific components in terms of downstream Sleep-EDF classification accuracy.
Method / Classification Sleep-EDF
BiTimelyGPT (Generative Pre-training) 90.41
    w/o BAAR 87.65
            w/o xPos (i.e., GPT-2) 84.25
BiTimelyGPT(w/o Pre-training) 86.33
5 Ablation Study
To evaluate the contributions of various components in BiTimelyGPT, we conducted ablation studies by selectively omitting bidirectionality and xPos embedding. We focused on the pre-training and fine-tuning on the Sleep-EDF dataset for classification. In Table 3, the xPos embedding increased performance by 3.40% by effectively encoding extrapolatable temporal patterns into time-series representations. The integration of the BAAR framework allows BiTimelyGPT to learn deep bidirectional representations, improving the downstream seq2vec tasks by 2.76%.
6 Probing Study
@@
We also explored the most effective choice of sequence features for downstream tasks (Table 4). Focusing on the last layer ($l=L$) and the last two layers ($l=\{L-1,L\}$), we examined four token combinations: [SOS] token alone, [EOS] token alone, both [SOS] and [EOS] tokens, and all tokens. In cases with multiple tokens, we applied average pooling for sequence representation. We observed that the [SOS] token from the last layer is sufficient to achieve the best Sleep-EDF classification, attributed to its aggregation of sequence information by backward attention.
Table 4: Probing study for the selection of sentence representations for downstream Sleep-EDF classification.
Sequence Represntation Sleep-EDF
[SOS] ($l=L$) 90.41
[EOS] ($l=L$) 89.66
[SOS]+ [EOS] ($l=L$) 90.11
All ($l=L$) 90.15
[SOS] ($l=\{L-1,L\}$) 88.63
[EOS] ($l=L$) 89.52
[SOS] + [EOS] ($l=\{L-1,L\}$) 89.22
All ($l=\{L-1,L\}$) 89.62
7 Conclusion and Future Work
@@
We present a BiTimelyGPT model, which leverages the recurrent attention and BAAR framework, for time-series representation learning and downstream discriminative tasks. Extensive visualization analyses underscore BiTimelyGPT’s three contributions: (1) the Next-Previous-Token Prediction pre-training effectively preserves the original distribution and data shapelets; (2) the BAAR framework alternates forward and backward retention across layers, learning deep bidirectional representations; (3) the full-rank forward and backward retention matrices enhance representation expressiveness. Empirical results demonstrate BiTimelyGPT’s superiority in the classification and regression on eight widely-used datasets. For future work, we would adapt BiTimelyGPT for out-of-distribution biosignals, further enhancing its utility in healthcare time-series. Additionally, we will broaden the applicability of BiTimelyGPT across multiple domains and data modalities, and experiment the benefits of BAAR framework with various autoregressive models.
@@ A.1 Denotations of Variables
$T$ Number of timesteps in a time-series $N$ Number of tokens
@@
$d$ The hidden dimension $h$ Number of heads
@@
[SOS] The start token of a sequence [EOS] The end token of a sequence
${\bm{Q}},{\bm{K}},{\bm{V}}\in\mathbb{R}^{N\times d}$ The query, key, value matrices ${\bm{W}}_{Q},{\bm{W}}_{K},{\bm{W}}_{V}\in\mathbb{R}^{d\times d}$ The projection matrices for ${\bm{Q}},{\bm{K}},{\bm{V}}$
$\overleftrightarrow{{\bm{A}}}\in\mathbb{R}^{N\times N}$ The bidirectional attention matrix $\overrightarrow{{\bm{A}}}\in\mathbb{R}^{N\times N}$ The forward attention matrix
${\bm{C}}\in\mathbb{R}^{N\times N}$ The lower triangular mask matrix ${\bm{S}}\in{\bm{R}}^{d\times d}$ The state variable for retention
@@
$\theta$ The rotary angle hyperparameter $\gamma$ The exponential decay hyperparameter
$\overrightarrow{{\bm{R}}}\in\mathbb{R}^{N\times N}$ The forward retention matrix $\overrightarrow{{\bm{D}}}\in\mathbb{R}^{N\times N}$ The lower triangular decay matrix
$\overleftarrow{{\bm{R}}}\in\mathbb{R}^{N\times N}$ The backward retention matrix $\overleftarrow{{\bm{D}}}\in\mathbb{R}^{N\times N}$ The upper triangular decay matrix
$\overleftrightarrow{{\bm{R}}}\in\mathbb{R}^{N\times N}$ The bidirectional retention matrix $\overleftrightarrow{{\bm{D}}}\in\mathbb{R}^{N\times N}$ The bidirectional decay matrix
@@ Appendix B Additional Information for Experiments
Dataset Task Features Timesteps Sequences Length Classes
Sleep-EDF (Kemp et al., 2000) Visualization 7 1.2B 400.7K 3,000 None
Sleep-EDF (sleep) (Kemp et al., 2000) Classification 1 586.4M 195.5K 3,000 5
Epilepsy (Andrzejak et al., 2001) Classification 1 2.0M 11.5K 178 2
PTB-XL (Alday et al., 2020) Classification 12 109.2M 21,837 5,000 5
EMG (Goldberger et al., 2000) Classification 1 306.0K 204 1,500 3
PPG-Dalia (Reiss et al., 2019a) Regression 4 16.6M 64,697 256 & $*{}^{*}$ None
IEEEPPG (Zhang et al., 2015) Regression 5 3.1M 3,096 1,000 None
RR (Pimentel et al., 2017) Regression 2 31.5M 7,870 4,000 None
HR (Pimentel et al., 2017) Regression 2 31.8M 7,949 4,000 None
SpO2 (Pimentel et al., 2017) Regression 2 31.8M 7,949 4,000 None

### v3

@@ 6.3 Qualitative analysis of attention matrix rankedness
Table 2: Comparison of BiTimelyGPT against various baselines for downstream seq2vec tasks. Metrics are reported as average (standard error) from a bootstrap evaluation of variance. The highest values per task (column) are highlighted in bold.
@@
BiTimelyGPT 90.4 (2.5) 93.0 (1.9) 87.5 (2.7) 95.4 (2.6) 26.1 (2.5) 2.81 (0.12) 8.42 (0.24) 4.23 (0.11)
TimelyGPT 89.2 (3.9) 92.8 (2.4) 86.5 (3.2) 95.9 (3.3) 26.2 (2.8) 2.78 (0.13) 8.53 (0.19) 4.26 (0.16)
TimelyBERT 82.2 (4.4) 85.6 (3.2) 81.0 (4.8) 84.5 (3.3) 31.0 (3.0) 4.22 (0.22) 12.86 (0.28) 4.82 (0.19)
CRT 90.1 (3.2) 91.1 (2.4) 87.8 (3.2) 94.6 (3.1) 26.5 (2.6) 2.96 (0.17) 9.02 (0.22) 4.48 (0.13)
PatchTST 89.6 (3.4) 91.3 (2.5) 83.4 (3.2) 95.2 (3.0) 26.1 (2.7) 2.89 (0.12) 9.46 (0.24) 4.45 (0.13)
AutoFormer 78.9 (3.6) 84.2 (2.5) 78.5 (3.5) 88.6 (3.7) 32.2 (2.9) 4.13 (0.15) 13.29 (0.28) 4.95 (0.13)
FedFormer 76.4 (3.6) 81.7 (2.7) 75.5 (3.5) 85.2 (2.7) 31.1 (2.5) 4.36 (0.21) 13.82 (0.32) 4.75 (0.15)
TimesNet 83.6 (3.5) 86.0 (2.7) 79.4 (3.2) 89.3 (3.2) 29.9 (2.7) 4.19 (0.18) 13.65 (0.27) 4.83 (0.16)
TST 88.8 (3.5) 88.0 (2.5) 81.9 (3.2) 94.2 (2.9) 26.8 (2.7) 3.47 (0.13) 12.63 (0.27) 4.95 (0.15)
GPT 84.3 (3.1) 88.7 (2.6) 83.1 (3.1) 92.6 (2.7) 28.6 (2.8) 3.71 (0.12) 11.09 (0.30) 4.64 (0.18)
GPT-BAAR 86.2 (3.5) 89.7 (2.8) 85.1 (3.1) 93.7 (3.3) 28.2 (2.8) 3.60 (0.19) 10.62 (0.27) 4.66 (0.13)
TS2Vec 86.2 (3.8) 88.3 (3.1) 82.7 (4.1) 93.8 (3.3) 27.9 (2.9) 3.53 (0.24) 11.56 (0.32) 4.60 (0.19)
TS-TCC 86.1 (3.9) 89.7 (2.9) 84.7 (3.9) 93.3 (3.8) 29.3 (3.2) 4.09 (0.23) 13.64 (0.35) 4.86 (0.19)
TF-C 86.6 (3.9) 87.5 (2.8) 82.7 (3.7) 93.8 (3.5) 28.5 (3.9) 4.38 (0.19) 14.15 (0.29) 4.87 (0.20)
LSTM 80.2 (3.9) 76.5 (3.2) 78.8 (3.5) 87.0 (2.9) 30.2 (3.2) 4.95 (0.22) 14.37 (0.41) 5.05 (0.23)
LSTM-BAAR 82.2 (3.5) 81.1 (2.9) 82.3 (3.1) 88.0 (3.0) 28.1 (2.8) 4.53 (0.14) 11.78 (0.28) 4.82 (0.19)
BiLSTM 81.6 (3.5) 78.3 (3.3) 80.7 (3.5) 86.8 (2.8) 29.4 (2.8) 4.71 (0.17) 13.50 (0.29) 4.97 (0.21)
ELMo 82.2 (3.5) 81.6 (3.0) 81.0 (3.1) 88.5 (2.9) 28.2 (3.0) 4.66 (0.15) 12.42 (0.24) 4.76 (0.15)
6.4 Quantitative results on discriminative fine-tuning tasks
@@
As shown in Table 2, BiTimelyGPT demonstrates the best performance in classification on the Sleep-EDF and Epilepsy datasets as well as regression on the IEEEPPG, HR and SpO2 datasets. These results highlighted the effectiveness of BiTimelyGPT’s bidirectionality, surpassing unidirectional TimelyGPT. Additionally, data preservation during pre-training and full-rank Retention matrices contribute considerably to BiTimelyGPT’s superior performance over other bidirectional transformers. CRT, employing a dropping-based pre-training, performs consistently well across datasets and achieves top results in the PTB-XL dataset. In contrast, masking-based PTMs typically lag behind due to the distribution shift. Among these masking-based PTMs, Autoformer, Fedformer, and TimesNet, which rely heavily on time decomposition and frequency-domain information, are especially affected by distribution shift. This suggested the advantage of dropping-based pre-training regime in avoiding distribution shift over the masking-based pre-training methods. Furthermore, both consistency-based PTMs and recurrent models show inferior performance in these tasks. The applications of the BAAR framework to GPT and LSTM result in improved performance, indicating its broad adaptability and effectiveness in enhancing various autoregressive models.
6.5 Classification of chronic disease from PopHR’s longitudinal records
BiTimelyGPT achieved superior performance over other time-series PTMs for classification of CHF, COPD, and Diabetes phenotypes (Table 3), which is attributable to its ability of utilizing bidirectional contexts. This proficiency in handling bidirectional information from irregularly sampled time series highlighted BiTimelyGPT’s potential in healthcare applications, where patients’ longitudinal records are often sporadic due to the irregular nature of patient visits. Notably, our BiTimelyGPT also surpasses other methods designed for irregularly sampled time series, leading with an average AUPRC of 60.5% and outperforming the next best mTAND (57.2%). Such results underscored the BiTimelyGPT’s generalizability, affirming its capability to model both continuous and irregularly sampled time series.
Table 3: Comparison of TimelyGPT and 9 baselines for multi-label classification (AUPRC %) on the three chronic diseases. Metrics are reported as average (standard error) from bootstrap evaluation of variance. Best values are highlighted in bold.
Model CHF COPD Diabetes
BiTimelyGPT 65.2 (2.2) 54.8 (1.6) 61.4 (1.4)
TimelyGPT 63.7 (3.3) 54.1 (2.8) 56.3 (2.4)
PatchTST 58.5 (3.5) 48.9 (2.9) 51.2 (2.2)
TST 59.3 (3.2) 49.7 (2.3) 52.0 (2.7)
CRT 54.2 (3.3) 44.6 (3.2) 46.7 (2.6)
TS-TCC 52.6 (4.2) 43.0 (3.9) 45.3 (3.7)
TS-T 50.5 (4.8) 40.9 (4.2) 43.1 (4.0)
mTAND 61.6 (1.9) 54.7 (1.6) 55.3 (1.4)
GRU-D 59.9 (2.6) 53.3 (2.1) 49.5 (2.0)
SeFT 54.8 (2.4) 45.2 (2.0) 47.5 (1.9)
RAINDROP 58.2 (3.1) 51.6 (2.5) 52.8 (2.6)
6.6 Ablation study
To evaluate different pre-training strategies while fixing the same neural architecture, we conducted ablation studies using the Sleep-EDF and PopHR datasets for classify sleep stages and CHF, respectively. We replaced the BAAR framework with two alternative pre-training strategies, namely (1) next-token prediction; (2) masking-based pre-training. We excluded the dropping-based method, which is tailored for encoder-decoder architectures. As shown in Table 4, the next-previous token prediction strategy achieves the highest performance, with 90.4% for Sleep-EDF and 65% for CHF. Replacing the BAAR framework with next-token prediction and masking-based pre-training resulted in a decline in performance for both datasets. Notably, the masking-based TimelyBERT underperforms compared to BiTimelyGPT even when our model is trained without pre-training, showing decreases of 4.1% for Sleep-EDF and 6.5% for CHF. This suggested that the Retention mechanism, which introduces decay in two directions, has less expressive representation power as discussed in Section 6.3. These findings highlighted the contribution of the BAAR framework and next-previous token Prediction in enhancing performance on discriminative tasks.
@@ 6.6 Ablation study
Table 4: Ablation results of BiTimelyGPT by assessing different pre-training strategies in terms of downstream classification tasks on Sleep-EDF and PopHR datasets.
Pre-training Strategies Sleep-EDF CHF (PopHR)
BiTimelyGPT with next-previous token prediction (Ours) 90.4 65.2
Replacing BAAR with next token prediction 87.7 63.2
Replacing BAAR with masking-based pre-training (i.e., TimelyBERT) 82.2 57.8
BiTimelyGPT without next-previous token prediction 86.3 64.3
6.7 Probing study
@@
We also explored the most effective choice of sequence features for downstream classification tasks using Sleep-EDF and PopHR datasets (Table 5). Focusing on the last layer ($l=L$) and the last two layers ($l=\{L-1,L\}$), we examined four token combinations: [SOS] token alone, [EOS] token alone, both [SOS] and [EOS] tokens, and average pooling of all tokens. We observed that the [SOS] token from the last layer achieves the best classification results, which can be attributed to its aggregation of sequence information by backward attention.
Table 5: Probing study on sentence representation for downstream classification tasks using Sleep-EDF and PopHR datasets, showing the best results with [SOS] in the final layer.
Sequence Represntation Sleep-EDF CHF (PopHR)
[SOS] ($l=L$) 90.4 65.2
[EOS] ($l=L$) 89.7 64.5
[SOS]+ [EOS] ($l=L$) 90.1 64.9
Average Pooling ($l=L$) 90.2 64.7
[SOS] ($l=\{L-1,L\}$) 88.6 64.4
[EOS] ($l=L$) 89.5 64.6
[SOS] + [EOS] ($l=\{L-1,L\}$) 89.2 64.6
Average Pooling ($l=\{L-1,L\}$) 89.6 64.4
7 Discussion
@@
We present a BiTimelyGPT model which leverages the recurrent attention and BAAR framework, for time-series representation learning and downstream discriminative tasks. Extensive visualization analyses underscore BiTimelyGPT’s three contributions: (1) the Next-Previous-Token Prediction pre-training effectively preserves the original distribution and data shapelets; (2) the BAAR framework alternates forward and backward Retention across layers, learning deep bidirectional representations; (3) the full-rank forward and backward Retention matrices enhance representation expressiveness. Empirical results demonstrate BiTimelyGPT’s superiority in classification and regression on eight datasets.
Limitations
This study concentrates on in-domain healthcare time-series data sharing similar biological patterns; however, its capacity to transfer learned representations to out-of-distribution biosignals remains untested. Future research should explore these out-of-distribution biosignals and data modalities to fully examine the model’s transferability across a wide range of healthcare data. Furthermore, while the BAAR framework’s effectiveness has been demonstrated with the Retention mechanism, its potential benefits for a variety of autoregressive models need further exploration.
8 Acknowledgements
Y.L. is supported by Canada Research Chair (Tier 2) in Machine Learning for Genomics and Healthcare (CRC-2021-00547), NSERC Discovery Grant (RGPIN-2016-05174), NOVA–FRQNT-NSERC (2023-NOVA-328677). Z.S. is supported by FRQ-NT Doctoral Training Scholarships.
@@ A.1 Denotations of Variables
$T$ Number of timesteps $N$ Number of tokens
@@
$d$ Hidden dimension $h$ Number of heads
@@
[SOS] Start token of a sequence [EOS] End token of a sequence
${\bm{Q}},{\bm{K}},{\bm{V}}\in\mathbb{R}^{N\times d}$ Query, key, value matrices ${\bm{W}}_{Q},{\bm{W}}_{K},{\bm{W}}_{V}\in\mathbb{R}^{d\times d}$ Projection matrices for ${\bm{Q}},{\bm{K}},{\bm{V}}$
$\overleftrightarrow{{\bm{A}}}\in\mathbb{R}^{N\times N}$ Bidirectional attention matrix $\overrightarrow{{\bm{A}}}\in\mathbb{R}^{N\times N}$ Forward attention matrix
${\bm{C}}\in\mathbb{R}^{N\times N}$ Lower triangular mask matrix ${\bm{S}}\in{\bm{R}}^{d\times d}$ State variable for retention
@@
$\theta$ Rotary angle hyperparameter $\gamma$ Exponential decay hyperparameter
$\overrightarrow{{\bm{R}}}\in\mathbb{R}^{N\times N}$ Forward retention matrix $\overrightarrow{{\bm{D}}}\in\mathbb{R}^{N\times N}$ Lower triangular decay matrix
$\overleftarrow{{\bm{R}}}\in\mathbb{R}^{N\times N}$ Backward retention matrix $\overleftarrow{{\bm{D}}}\in\mathbb{R}^{N\times N}$ Upper triangular decay matrix
$\overleftrightarrow{{\bm{R}}}\in\mathbb{R}^{N\times N}$ Bidirectional retention matrix $\overleftrightarrow{{\bm{D}}}\in\mathbb{R}^{N\times N}$ Bidirectional decay matrix